Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 117     Entry time: 31 Mar 2023, 05:10
Author: Benjamin, Yi-Ping, Ru-Pan, Andreas 
Group: SCS 
Subject: Late Shift Summary 

Achievements

  • Optimized BOZ for Co and Fe
  • Estabilished Stable liquid jet flow with 30 mM Fe(bpy)3

Problems

  • Jet froze several times
  • Chamber warmed up earlier in the day. This led to a "dusting" of CNCbl on many surfaces and in the large turbo. We cleaned the turbo with ethanol.

Plan for Night Shift

  • Continue measuring stitic XAS with BOZ scheme.
ELOG V3.1.4-7c3fd00